# Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA

### v6

@@ Abstract
Continual learning (CL) aims to accumulate knowledge from sequential tasks without catastrophic forgetting. Vision–language models like CLIP, with strong generalization, are widely used for CL. Existing methods often adapt isolated PTM components, adding inference complexity and limiting PTM improvement, or rely on replay, stored information, or assumptions, incurring high costs and limited applicability. To advance models as continual learners, we explore CL via natural, efficient PTM updates instead of complex task-specific additions. We thus study continual low-rank learning and systematically analyze how LoRA ranks and placements affect learning and forgetting. We find that a relatively higher-rank LoRA improves task learning (i.e., plasticity) but increases forgetting, while a relatively lower-rank LoRA reduces forgetting (i.e., stability) but limits adaptation. Crucially, we find a plasticity–stability balance tied to rank across parameters and tasks, with moderately small ranks maximizing CL benefits. Motivated by this, we propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which continually updates PTMs with LoRA adapters for all parameters of adaptively optimized rank number. While the new-task objective drives learning, CoDyRA adaptively minimizes ranks with sparsity-promoting regularization to reduce interference and forgetting, achieving a plasticity–stability balance tailored to different parameters and tasks. Despite allowing updates to all parameters, the adaptively selected and minimized LoRA ranks help keep the updated model closer to its previous state, while still enabling comprehensive and effective learning of new tasks. CoDyRA enables efficient CL as a sequence of LoRA-based tasks without storing past data, task information, or relying on assumptions. It preserves the original model architecture and deployment pipeline, adding no inference overhead. Extensive experiments show CoDyRA improves new representations while retaining old knowledge, achieving state-of-the-art results. Code is available.
@@ 1 Introduction
Continual learning (CL) (Hadsell et al., 2020; De Lange et al., 2021; Wang et al., 2024b) focuses on incrementally learning from new data streams without catastrophic forgetting (Nguyen et al., 2019; McCloskey & Cohen, 1989) or the need to retrain on all previously seen data. With the rise of large-scale pre-trained models (PTMs), recent studies increasingly perform CL in this context (Zhou et al., 2024; McDonnell et al., 2024; Wang et al., 2024a; Jha et al., 2024). Many PTM-based CL approaches primarily focus on continual adaptation for downstream tasks (e.g., class-incremental learning, CIL (Smith et al., 2023; Wortsman et al., 2022)), with an emphasis on downstream tasks. While task-specific and downstream-oriented designs (Zhou et al., 2024; McDonnell et al., 2024; Xu et al., 2024; Yu et al., 2024) show promising results on particular CL tasks, PTMs remain largely isolated from the CL process, mitigating forgetting but undermining interaction with new tasks and thus the potential for continual improvement. To advance toward models that function as continual learners, we consider CL with PTMs that enable flexible and efficient updates, preserving or even enhancing generalizable representations while acquiring new knowledge.
@@
Leveraging the strong generalization and flexibility of jointly learned vision–text embeddings, CLIP as a vision–language model has been widely adopted in CL (Jha et al., 2024; Zheng et al., 2023; Wang et al., 2023; Zhou et al., 2023), including continual updating (Garg et al., 2023) and multi-domain CL (Zheng et al., 2023). Beyond downstream performance, the retention of knowledge in PTMs is also evaluated. Some approaches, such as ZSCL (Zheng et al., 2023), aim to continually update PTM, i.e., CLIP, through full fine-tuning while mitigating forgetting via memory replay (Li & Hoiem, 2017; Wortsman et al., 2022; Rebuffi et al., 2017b) using additional reference data. However, this leads to high training costs and still struggles with catastrophic forgetting (Fig. 1(a)). To directly control the forgetting, some methods (Yu et al., 2024; Xu et al., 2024) introduce task-specific components isolated from the frozen PTMs, requiring task/domain prediction or gating at inference. These methods can deliver strong results with task-specific gating and branching; however, by isolating tasks and relying heavily on per-task modules and accurate task prediction, they become restricted to certain domains and limited in both general CL applicability and inference efficiency. Moreover, with restricted interaction between PTMs and CL process (e.g., isolation) limits PTMs’ generalization improvement potential (Fig. 1(b)).
@@
Figure 1: (a) Continual full fine-tuning leverages reference data to reduce forgetting of pre-trained knowledge (e.g., zero-shot performance (Zheng et al., 2023)) but remains costly and prone to forgetting. (b) Task-specific modular methods (e.g., (Yu et al., 2024; Xu et al., 2024)) add isolated components requiring domain prediction or gating, increasing complexity and limiting adaptability and generalization. (c) Our approach enables efficient, universal continual PTM updating without reference data or task-specific designs, retaining PTM knowledge and improving performance on unseen data.
@@
Straightforward adaptation of PTMs can be achieved through direct model updating through full fine-tuning (FT) or parameter-efficient fine-tuning (PEFT) methods (Zhou et al., 2022c; b), such as Low-Rank Adaptation (LoRA) (Hu et al., 2021). This raises a fundamental question: instead of relying on complex additions or task-specific components, can CL be performed by leveraging these natural PTM update mechanisms? Our work investigates how to mitigate catastrophic forgetting within these standard frameworks without imposing extensive assumptions. We propose an efficient CL approach for CLIP that enables the model to be updated naturally across a stream of tasks, while retaining foundational pre-trained knowledge. We build upon LoRA for efficiency. While there have been some LoRA-based CL methods (Liang & Li, 2024; Wu et al., 2025; Zhu et al., 2025), they rely on specific regularization (e.g., orthogonal property) for specific settings.
@@
The analysis results motivate us to propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which continually updates the PTM using LoRA adapters with dynamically and adaptively optimized ranks. CoDyRA performs new task learning while jointly minimizing the rank of each LoRA adaptively. Rank selection is guided by sparsity-promoting regularization and learnable importance weights, which dynamically adjust the contribution of each rank during training. By updating with LoRA adapters of minimized ranks, CoDyRA inherently biases models toward remaining closer to their previous state, thereby alleviating forgetting in principle without task-specific assumptions.
Our contributions are fourfold: (1) We study PTM-based CL with natural and efficient continual update of the model, not relying on complex task-specific additions, without storing past data, task information, or relying on assumptions. (2) We comprehensively analyze learning–forgetting trade-offs across LoRA ranks and placements, identifying a plasticity–stability balance at moderately small ranks and deriving design strategies. This study guides our CL approach with regularized forgetting and provides evidence for future research. (3) We propose CoDyRA, which adaptively minimizes LoRA ranks via sparsity-promoting regularization. It handles different parameters and tasks, yielding regularized updates that reduce interference and keep the model closer to its previous state with less forgetting. (4) Extensive experiments across multiple CL benchmarks validate our method’s superior performance in both downstream learning and mitigating forgetting, achieving a balanced trade-off between plasticity and stability.
@@ 2 Related Work
Continual Learning. Continual learning enables models to sequentially learn new tasks while retaining previously acquired knowledge. Experience replay (ER) methods (Luo et al., 2023; Aljundi et al., 2019b; Chaudhry et al., 2018a; Liu et al., 2020; Chaudhry et al., 2018b; Yan et al., 2022; 2021) store subsets of past data to refresh the model on prior tasks during new training. Parameter regularization approaches (Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017; Aljundi et al., 2019a; Jha et al., 2023) constrain updates to important weights to preserve past knowledge. Meanwhile, dynamic networks (Wang et al., 2022a; b; Zhou et al., 2022a; Wang et al., 2024a; McDonnell et al., 2024; Liang & Li, 2024; Wang et al., 2022e; d; Smith et al., 2023; Wang et al., 2022c) adjust their architecture on the fly, balancing the acquisition of new information with the retention of old knowledge.
Continual Learning of CLIP. Continual learning of CLIP (Jha et al., 2024; Zhang et al., 2024b) aims to enable PTMs to sequentially learn across diverse domains while retaining their pre-trained generalization capabilities for previously seen tasks. (Garg et al., 2023) continually adapt CLIP to temporal distribution shifts via timestamped data updates. The approach introduced in ZSCL (Zheng et al., 2023) addresses the challenge of preserving zero-shot capabilities while adapting to new tasks, mitigating the risk of catastrophic forgetting. Subsequent works (Yu et al., 2024; Tang et al., 2025; Xu et al., 2024; Huang et al., 2025) have focused primarily on continual adaptation to downstream tasks while leveraging pre-trained predictions for unseen data. Recently, (Xu et al., 2024) proposed a more challenging setting, where the labels from different domains are mixed during testing.
Low-Rank Adaptation. Low-Rank Adaptation (LoRA) (Hu et al., 2021) has gained popularity as a parameter-efficient method for fine-tuning large pre-trained models. Similar to LoRA, (Wu et al., 2024b) learns a small set of task‐specific representation vectors at each layer. Building on LoRA’s framework, recent works enhance LoRA by reformulating parameter updates, focusing on initialization based on the pre-trained weights (Zhang et al., 2024a; Meng et al., 2024), treating each rank independently (Ding et al., 2023; Zhang et al., 2023; Liu et al., 2024), or employing a mixture of subspaces (Wu et al., 2024a). These advancements highlight LoRA’s effectiveness in efficiently fine-tuning large models while preserving generalization.
@@ 3.1 Preliminaries
Transformers. A typical transformer model is composed of stacked blocks, each containing two main submodules: an multi-head attention (MHA) module (Attn) and a multilayer perceptron (MLP). The attention module at the $l$-th layer is defined as: $\text{Attn}(x)=\text{MHA}(x\mathbf{W}^{\text{Q}}_{l},x\mathbf{W}^{\text{K}}_{l},x\mathbf{W}^{\text{V}}_{l})\mathbf{W}^{\text{O}}_{l}.$ The input sequence is denoted as $x\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the hidden dimension. The matrices $\mathbf{W}^{\text{Q}}_{l},\mathbf{W}^{\text{K}}_{l},\mathbf{W}^{\text{V}}_{l},\mathbf{W}^{\text{O}}_{l}\in\mathbb{R}^{d\times d}$ correspond to the Query, Key, Value, and Output projection matrices, respectively. The MLP in each transformer block consists of two linear layers separated by an activation function $\text{Act}(\cdot)$, such as ReLU or GELU (Hendrycks & Gimpel, 2016). The MLP module at the $l$-th transformer layer is defined as: $\text{MLP}(x)=\text{Act}(x\mathbf{W}^{\text{FC}}_{l}+b^{\text{FC}}_{l})\mathbf{W}^{\text{Proj}}_{l}+b^{\text{Proj}}_{l},$ where $\mathbf{W}^{\text{FC}}_{l}\in\mathbb{R}^{d\times d_{m}}$ and $\mathbf{W}^{\text{Proj}}_{l}\in\mathbb{R}^{d_{m}\times d}$ are the weight matrices, with $d_{m}$ being the hidden dimension.
CLIP model. In CLIP (Radford et al., 2021; Ilharco et al., 2021; Fang et al., 2024; Li et al., 2022), both the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are transformers that learn a joint vision–language embedding. Given an image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, the vision encoder produces features ${\mathbf{z}}_{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$. A class label $y$ is embedded into a prompt such as “a photo of a [CLS]” to form text input $\mathbf{t}$}, which the text encoder maps to ${\mathbf{z}}_{\text{T}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$. At inference, the probability of classifying $\mathbf{x}$} into class $y_{i}\in\{1,\dots,C\}$ is: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}_{\text{V}},{\mathbf{z}}_{\text{T}}^{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}_{\text{V}},{\mathbf{z}}_{\text{T}}^{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity and $\tau$ is a temperature parameter.
@@
CL with CLIP. We study continual learning of CLIP with multi-domain data, e.g., multi-domain task-incremental learning (MTIL) (Zheng et al., 2023; Yu et al., 2024) and cross-domain task-agnostic incremental learning (X-TAIL) (Xu et al., 2024). In MTIL, the model sequentially learns from $T$ tasks, where each task $t$ has dataset $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, with input image ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$, label $y^{t}_{i}\in\mathcal{C}^{t}$, and $M^{t}$ classes in $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$. In X-TAIL, the test-time category set includes both seen and unseen domains, $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$, where $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ contains all classes from prior tasks and $\mathcal{C}_{\text{unseen}}$ denotes novel classes never encountered.

### v7

# Adaptive Rank, Reduced Forgetting: Continual Learning with Dynamic Rank-Selective LoRA

@@ Abstract
Continual learning (CL) aims to accumulate knowledge from sequential tasks without catastrophic forgetting. Many CL methods for pre-trained models (PTMs) achieve strong performance but rely on sophisticated architectures or additional assumptions. They adapt isolated modules—adding inference complexity and limiting PTM improvement—or depend on replay, stored data, or assumptions, incurring extra cost and reduced applicability. To advance usage, we explore natural, efficient PTM CL instead of complex task-specific additions. We study continual low-rank learning and systematically analyze how LoRA ranks and placements affect learning and forgetting. We find that a relatively higher-rank LoRA improves task learning (i.e., plasticity) but increases forgetting, while lower-rank LoRA reduces forgetting (i.e., stability) but limits adaptation. Crucially, we observe a plasticity–stability balance tied to ranks across parameters and tasks, with moderately small ranks maximizing CL benefits. Motivated by this, we propose Continual Dynamic Rank-Selective LoRA (CoDyRA), continually updating PTMs with LoRA of adaptively optimized ranks. While the new-task objective drives learning, CoDyRA adaptively minimizes ranks with sparsity-promoting regularization to reduce interference and forgetting, achieving plasticity–stability balance adaptively for different parameters and tasks. It preserves the original model architecture and serving pipeline, adding no inference overhead. Extensive experiments on both CLIP and LLMs show CoDyRA improves new-task learning with strong knowledge retention, delivering state-of-the-art results. Code is available.
@@ 1 Introduction
Continual learning (CL) [23, 11, 78] focuses on incrementally learning from new data streams without catastrophic forgetting [56, 51] or the need to retrain on all previously seen data. With the rise of pre-trained models (PTMs), recent studies increasingly perform CL in this context [101, 52, 75, 31]. Following the context of class-incremental learning (CIL) tasks [68, 85], many PTM-based CL approaches primarily focus on or are evaluated through continual adaptation to downstream tasks. While task-specific and downstream-oriented designs [101, 52, 90, 93] show promising results on particular CL tasks, PTMs remain largely isolated from the CL process, mitigating forgetting but undermining interaction with new tasks and thus the potential for continual improvement. To advance toward models that function as continual learners, we consider CL with PTMs that enable flexible and efficient updates, preserving or even enhancing generalizable representations while acquiring new knowledge.
@@
Leveraging the strong generalization and flexibility of jointly learned vision–text embeddings, CLIP as a vision–language model has been widely adopted in CL [31, 99, 79, 103], including continual updating [21] and multi-domain CL [99]. Beyond downstream performance, the retention of knowledge in PTMs is also evaluated. Some approaches, such as ZSCL [99], aim to continually update the pretrained CLIP through full fine-tuning while mitigating forgetting via memory replay [38, 85, 64] using additional reference data. However, this leads to high training costs and still struggles with catastrophic forgetting (Fig. 1(a)). To directly control the forgetting, some methods [93, 90] introduce task-specific components isolated from the frozen PTMs, requiring task/domain prediction or gating for routing the specific components at inference. These methods can deliver strong results with task-specific gating and branching; however, by isolating tasks and relying heavily on per-task modules and accurate task prediction, they become restricted to certain domains and limited in both general CL applicability and inference efficiency. Moreover, the restricted interaction between PTMs and CL process (e.g., isolation) limits PTMs’ generalization improvement potential (Fig. 1(b)).
@@
Figure 1: (a) Continual full fine-tuning leverages reference data to reduce forgetting of pre-trained knowledge (e.g., zero-shot performance [99]) but remains costly and prone to forgetting. (b) Task-specific modular methods (e.g., [93, 90]) add isolated components requiring domain prediction or gating, increasing complexity and limiting adaptability and generalization. (c) Our approach enables efficient, universal continual PTM updating without reference data or task-specific designs, retaining PTM knowledge and improving performance on unseen data.
@@
General-purpose PTM updating can be achieved through direct model updating via straightforward full fine-tuning (FT) or parameter-efficient fine-tuning (PEFT) methods [105, 104], such as Low-Rank Adaptation (LoRA) [27]. This raises a fundamental question: instead of relying on complex additions or task-specific components, can CL be performed by leveraging these natural PTM update mechanisms? Our work investigates how to mitigate catastrophic forgetting within these standard frameworks without imposing additional assumptions. We propose an efficient CL approach for CLIP that enables the model to be updated naturally across a stream of tasks, while retaining foundational pre-trained knowledge. We build upon LoRA for efficiency. There have been some LoRA-based CL methods [40, 87, 106], which rely on specific regularization (e.g., orthogonal property) for specific settings.
@@
The analysis results motivate us to propose Continual Dynamic Rank-Selective LoRA (CoDyRA), which continually updates the PTM using LoRA adapters with dynamically and adaptively optimized ranks. CoDyRA performs new task learning while jointly minimizing the rank of each LoRA adaptively. Rank selection is guided by sparsity-promoting regularization and learnable importance weights, which dynamically adjust the contribution of each rank and decide a balanced capacity in training. By updating with LoRA adapters of minimized ranks, CoDyRA inherently regularizes models toward remaining closer to their previous state, thereby alleviating forgetting in principle without task-specific assumptions.
Our contributions are fourfold: (1) We study PTM-based CL with natural and efficient continual update of the model, not relying on complex task-specific additions, without storing past data, task information, or relying on assumptions. (2) We comprehensively analyze learning–forgetting trade-offs across LoRA ranks and placements, identifying a plasticity–stability balance at moderately small ranks and deriving design strategies. This study guides our CL approach with regularized forgetting and provides evidence for future research. (3) We propose CoDyRA, which adaptively minimizes LoRA ranks via sparsity-promoting regularization. It handles different parameters and tasks, yielding regularized updates that reduce interference and keep the model closer to its previous state with less forgetting. (4) Extensive experiments across multiple CL benchmarks demonstrate superior downstream performance and reduced forgetting, achieving a strong plasticity–stability balance. We further extend CoDyRA to LLMs beyond CLIP, validating its effectiveness on broader applications and pre-trained models.
@@ 2 Related Work
Continual Learning. Continual learning enables models to sequentially learn new tasks while retaining previously acquired knowledge. Experience replay (ER) methods [48, 3, 8, 42, 9, 91, 92] store subsets of past data to refresh the model on prior tasks during new training. Parameter regularization approaches [33, 1, 94, 2, 32, 39] constrain updates to important weights to preserve past knowledge. Dynamic networks [73, 74, 102, 75, 52, 40, 84, 83, 68, 82, 65] adjust their architecture on the fly, balancing the acquisition of new information with the retention of old knowledge.
Continual Learning of CLIP. CL of CLIP [31, 97] aims to enable PTMs to sequentially learn across diverse domains while retaining their pre-trained generalization capabilities for previously seen tasks. [21] continually adapt CLIP to temporal distribution shifts via timestamped data updates. ZSCL [99] was proposed to address catastrophic forgetting while preserving zero-shot capabilities during task adaptation. Subsequent works [93, 69, 90, 29, 41] have focused primarily on continual adaptation to downstream tasks while leveraging pre-trained predictions for unseen data. Recently, a more challenging setting was proposed, where the labels from different domains are mixed during testing [90].
Low-Rank Adaptation. Low-Rank Adaptation (LoRA) [27] has gained popularity as a parameter-efficient method for fine-tuning large pre-trained models. Similar to LoRA, [88] learns a small set of task‐specific representation vectors at each layer. Building on LoRA’s framework, recent works enhance LoRA by reformulating parameter updates, focusing on initialization based on the pre-trained weights [95, 53], treating each rank independently [14, 96, 43], or employing a mixture of subspaces [86]. LoRA has been shown to be effective even with small ranks [5, 66], while our work investigates how to perform low-rank learning in CL and how to dynamically and adaptively decide capacity. Owing to their effectiveness and efficiency, LoRA-based methods have been extended to CL of LLMs, often relying on specialized regularization or architectural designs [80, 61, 60]. For instance, O-LoRA [80] mitigates interference via orthogonal regularization; TreeLoRA [60] routes inputs to task-specific adapters using a gradient-similarity tree with bandit search.
@@ 3.1 Preliminaries
Transformers. A typical transformer model is composed of stacked blocks, each containing two main submodules: an multi-head attention (MHA) module (Attn) and a multilayer perceptron (MLP). The attention module at the $l$-th layer is defined as: $\text{Attn}(x)=\text{MHA}(x\mathbf{W}^{\text{Q}}_{l},x\mathbf{W}^{\text{K}}_{l},x\mathbf{W}^{\text{V}}_{l})\mathbf{W}^{\text{O}}_{l}.$ The input sequence is denoted as $x\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the hidden dimension. The matrices $\mathbf{W}^{\text{Q}}_{l},\mathbf{W}^{\text{K}}_{l},\mathbf{W}^{\text{V}}_{l},\mathbf{W}^{\text{O}}_{l}\in\mathbb{R}^{d\times d}$ correspond to the Query, Key, Value, and Output projection matrices, respectively. The MLP in each transformer block consists of two linear layers separated by an activation function $\text{Act}(\cdot)$, such as ReLU or GELU [26]. The MLP module at the $l$-th transformer layer is defined as: $\text{MLP}(x)=\text{Act}(x\mathbf{W}^{\text{FC}}_{l}+b^{\text{FC}}_{l})\mathbf{W}^{\text{Proj}}_{l}+b^{\text{Proj}}_{l},$ where $\mathbf{W}^{\text{FC}}_{l}\in\mathbb{R}^{d\times d_{m}}$ and $\mathbf{W}^{\text{Proj}}_{l}\in\mathbb{R}^{d_{m}\times d}$ are the weight matrices, with $d_{m}$ being the hidden dimension.
CLIP model. In CLIP [62, 30, 17], both the image encoder $f_{\theta}$ and text encoder $g_{\psi}$ are transformers that learn a joint vision–language embedding. Given an image ${\mathbf{x}}\in\mathbb{R}^{H\times W\times C}$, the vision encoder produces features ${\mathbf{z}}_{\text{V}}=f_{\theta}({\mathbf{x}})\in\mathbb{R}^{D}$. A class label $y$ is embedded into a prompt such as “a photo of a [CLS]” to form text input $\mathbf{t}$}, which the text encoder maps to ${\mathbf{z}}_{\text{T}}=g_{\psi}({\mathbf{t}})\in\mathbb{R}^{D}$. At inference, the probability of classifying $\mathbf{x}$} into class $y_{i}\in\{1,\dots,C\}$ is: $p(y_{i}|{\mathbf{x}})=\frac{\exp(\mathtt{sim}({\mathbf{z}}_{\text{V}},{\mathbf{z}}_{\text{T}}^{y_{i}})/\tau)}{\sum_{c=1}^{C}\exp(\mathtt{sim}({\mathbf{z}}_{\text{V}},{\mathbf{z}}_{\text{T}}^{y_{c}})/\tau)},$ where $\mathtt{sim}(\cdot)$ is cosine similarity and $\tau$ is a temperature parameter.
CL with CLIP. We study continual learning of CLIP with multi-domain data, e.g., multi-domain task-incremental learning (MTIL) [99, 93] and cross-domain task-agnostic incremental learning (X-TAIL) [90]. In MTIL, the model sequentially learns from $T$ tasks, where each task $t$ has dataset $\mathcal{D}^{t}=\{({\mathbf{x}}^{t}_{i},y^{t}_{i})\}_{i=1}^{N^{t}}$, with input image ${\mathbf{x}}^{t}_{i}\in\mathbb{R}^{H\times W\times C}$, label $y^{t}_{i}\in\mathcal{C}^{t}$, and $M^{t}$ classes in $\mathcal{C}^{t}=\{y^{t}_{j}\}_{j=1}^{M^{t}}$. In X-TAIL, the test-time category set includes both seen and unseen domains, $\mathcal{C}=\mathcal{C}_{\text{seen}}\cup\mathcal{C}_{\text{unseen}}$, where $\mathcal{C}_{\text{seen}}=\bigcup_{i=1}^{n}\mathcal{C}_{i}$ contains all classes from prior tasks and $\mathcal{C}_{\text{unseen}}$ denotes novel classes never encountered.